OpenStudy (anonymous):

Let f(x)=x^2+5. At some point on the graph f(x), the tangent line is parallel to 6x-100. Write down the equation of this line.

OpenStudy (aum):

What is the slope of the line 6x - 100 ?

OpenStudy (anonymous):

6

OpenStudy (aum):

What is the slope of f(x) in terms of x?

OpenStudy (anonymous):

are you asking about the slope of f(x)=x^2+5

OpenStudy (anonymous):

?

OpenStudy (aum):

The first derivative gives the slope of f(x). \[ f(x) = x^2 + 5 \\ f'(x) = 2x \]

OpenStudy (aum):

If two lines are parallel, their slopes must be the same. Therefore, 2x = 6 x = 3. when x = 3, f(x) = 3^2 + 5 = 9 + 5 = 14 Therefore, a tangent to f(x) drawn at the point (3, 14) will be parallel to the line 6x - 100. Equation of the tangent is: y = mx + b y = 6x + b (3,14) is a point on the line: 14 = 6*3 + b b = -4 y = 6x - 4 is the equation of the tangent.

OpenStudy (anonymous):

Thank you !!!

OpenStudy (aum):

You are welcome.